20Januari,,en,De deuren van de Gardeners kamer hebben eindelijk afgelopen en zijn vervangen door een paar Aluminium deuren,,en,nieuwe lampen worden geleverd en gemonteerd rond het midden zwembad aangezien dit heeft bewezen extra verlichting nodig hebben en op elk gewenst moment werden de Pool verlichting werkte niet,,en,Picon is afgeleverd en wordt momenteel verspreid over de tuinen,,en,Alexis heeft gepland in de Palm Tree snoeien en zal zo snel mogelijk worden begonnen,,en,De schimmel / witte vlieg is momenteel onder controle, maar we zullen in geval van verdere besmetting waakzaam blijven,,en,In de komende weken zullen we doorgaan met de straat licht kabelvervanger,,en 2014

We have installed 2 new amplifier houses. One on the Ts and Vs and one adjacent to No.12. This will improve TV reception.

As we have previously mentioned the insulation cover on the street lights is breaking down and the earth spikes are corroded owing to the salt in the soil. We have decided to install new street light cables and spikes in the areas we have not already covered. Although Industria have not replied to our request for information we are working to the latest regulations. This will alleviate the problems with the lights that some areas are currently experiencing.

The Heat Pump while the initial fault was Electrical and dealt with, a minor leak has also been found in the Gas. This is currently being dealt with. The Pump is under five years old so is still under warranty.

The annual pruning of the Palm Trees is underway and will continue for several weeks.

Brian

 

16Januari,,en,De deuren van de Gardeners kamer hebben eindelijk afgelopen en zijn vervangen door een paar Aluminium deuren,,en,nieuwe lampen worden geleverd en gemonteerd rond het midden zwembad aangezien dit heeft bewezen extra verlichting nodig hebben en op elk gewenst moment werden de Pool verlichting werkte niet,,en,Picon is afgeleverd en wordt momenteel verspreid over de tuinen,,en,Alexis heeft gepland in de Palm Tree snoeien en zal zo snel mogelijk worden begonnen,,en,De schimmel / witte vlieg is momenteel onder controle, maar we zullen in geval van verdere besmetting waakzaam blijven,,en,In de komende weken zullen we doorgaan met de straat licht kabelvervanger,,en 2014

 

Both pool Heat Pumps experienced problems last week. The Pool at the Bar had a gas leak. This has been repaired at a cost of 256,80 euro. This includes labour refrigerant and Gas.

The other Heat Pump is an electrical fault. The display unit face had a been taken off by persons unknown and condensation had affected the computerized board. This is currently being dealt with.

The center road wall repairs have been completed on one side by Todao and should result in a lot less maintenance in the future.

 

1st December

We are continuing to install new cable for street lighting.

The wall to the centre road is being refurbished by Todao and is looking good.

As many of the bases to the street lights are corroding making changing bulbs and globes on a ladder dangerous we are currently trialing new lower lights which also eliminate light pollution.

Existing walkway manhole covers being improved and repaired for safer walking.

The Bar refurbishment is almost complete and should reopen in the next few weeks .

An extra telephone line is being installed to assist the wi fi system to half the loading from the TV room, as we are nearing capacity for one line to the area

Brian

22nd oktober,,en,We zijn nog steeds te maken met de besmetting op de Hibiscus,,en,De deskundigen advies is om elke vijftien dagen spuiten en twee dagen later af te wassen de planten,,en,Het duurt enkele dagen om deze taak te voltooien en is uiteraard zo arbeidsintensief dat we niet kan bijhouden met andere werken in de tuin,,en,De enige manier om een ​​succesvolle oplossing voor het probleem te bereiken is om een ​​vierde persoon in dienst op tijdelijke basis voor het uitvoeren van de benodigde werkzaamheden aan de Hibiscus en de tuinders vrij te geven voor hun normale werk,,en,Wanneer de eigenaars hebben gevraagd we hebben verwijderd Hibiscus die worden beïnvloed door de besmetting,,en,We hebben onlangs vervangen oude coaxkabel in,,en,gebieden waar ze degradatie als gevolg had geleden om de blootstelling aan de zon,,en,De slapende agenten lijken te hebben van de gewenste invloed,,en 2013

Since reporting on the street lighting issues in the Autumn newsletter the situation has unfortunately deteriorated further. We have found additional cables that have failed because of damage from long term water immersion in the macrons ( the underground pipes which carry the cables). This has caused problems with the lights down the main road to the gate and along the side road behind 200.

The electrician along with the help of Hedley (203) is managing temporary solutions, but these cables will also need replacing as part of the refurbishment of the lighting system.

Brian

19th October 2013

Fungus /White Fly

We have made contact with an expert regarding the current problem with the Plants. We are using the correct chemicals for the problem. The only additional treatment is to hose down the plants after treating with chemicals, which we will now implement.

We are currently repairing the Fence damaged near Casa J by persons unknown. This Fence is generally in a poor state and will require further work in the near future.

We are continuing to upgrade street light cable and have now completed another section. This work will continue while the problem persists.

Brian

2nd oktober,,en,We zijn nog steeds te maken met de besmetting op de Hibiscus,,en,De deskundigen advies is om elke vijftien dagen spuiten en twee dagen later af te wassen de planten,,en,Het duurt enkele dagen om deze taak te voltooien en is uiteraard zo arbeidsintensief dat we niet kan bijhouden met andere werken in de tuin,,en,De enige manier om een ​​succesvolle oplossing voor het probleem te bereiken is om een ​​vierde persoon in dienst op tijdelijke basis voor het uitvoeren van de benodigde werkzaamheden aan de Hibiscus en de tuinders vrij te geven voor hun normale werk,,en,Wanneer de eigenaars hebben gevraagd we hebben verwijderd Hibiscus die worden beïnvloed door de besmetting,,en,We hebben onlangs vervangen oude coaxkabel in,,en,gebieden waar ze degradatie als gevolg had geleden om de blootstelling aan de zon,,en,De slapende agenten lijken te hebben van de gewenste invloed,,en 2013

Hi

We have now repainted the Promenade Gate area, including the walls and gate, after the keypad damage. The Gate Closer has unfortunately been vandalised again and will be replaced with a stronger closure.

We are spraying the plants and bushes again for fungus as the Island still has the fungal problem.

We have now spoken with Canal Gestion, the new water company, and hopefully improved the water pressure to the gardens. We appreciate the positive messages we have received regarding the way plants are growing with the new system.

All the Team are now back at work after the Summer Holidays.

Brian

 

11th September 2013

Virtually all suspect cabling for the lights has been replaced. What has become apparent during this operation is that the insulating cable has broken down and has caused the problems. Additionally, it is been impossible to use the existing ducts in places. This has made it a larger project than we envisaged.

Miguel is currently on holiday and by September end the Team will have used all their holiday entitlement.

The new water company, initially insisted on installing extra filters to the system which effected our pressure. They have since agreed to remove them and we should now have improved pressure.

Brian

13th August 2013

Further to the last blog about the issues we have been having with the complex lights, it is becoming apparent that the oldest area of the complex has cabling past its sell by date. The replacement of Cables is proving more extensive than we envisaged at the start. It is more the cable insulation on the outside that’s broken down rather than the cable itself.

Hedley Smith 203, has been a great help with this problem and liaises with the electrician.

Brian

5th August 2013

The rewiring of the Lights has now been completed in the affected areas and all complex lights are now working. This has been a difficult re-cabling project. Now the earth stakes are being installed.

For a period of 2 weeks from Tuesday 6th August, reception will have a limited opening time from 9.00am to 1.00pm.

 

Brian

26th July 2013

Er wordt verder gewerkt aan het oplossen van de problemen met de straatverlichting en de corroderende kabels. Een aantal van de ondergrondse leidingen waar de straatlantaarnkabels doorheen lopen, zijn nu volledig ingestort. We moeten nieuwe kabels omleiden door nieuwe kabelgoten en op een of twee plekken de weg uitgraven. Uiteraard moet er scherp op de kosten worden gelet, maar dit werk is onvermijdelijk en zal deel uitmaken van het werk dat Industria ons uiteindelijk zal vragen uit te voeren.

 

Brian